# Zeroth-Order Constrained Optimization from a Control Perspective via Feedback Linearization

### v1

@@ Appendix C Proof of Theorem 2
$\|\lambda-\lambda^{\star}\|=\|\lambda(1)-\lambda(0)\|\leq\left(\frac{4096n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{4}}+\frac{64L_{f}}{\underline{L}_{h}^{2}}\right)nRr_{2}^{2}.$
Thus
$\quad\|J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}(\lambda_{t}^{\star}-\lambda_{t})\|\leq n\overline{L}_{h}^{2}\|\lambda_{t}^{\star}-\lambda_{t}\|$
$\leq n^{2}\overline{L}_{h}^{2}R\left(\frac{4096n\overline{L}_{h}(L_{f}\overline{L}_{h}+\|K\|H)}{\underline{L}_{h}^{4}}+\frac{64L_{f}}{\underline{L}_{h}^{2}}\right)r_{2}^{2}$
Appendix D Bounding ${\lambda_{\min}}(J_{h}(x_{t})\widetilde{J}_{h}(x_{t}))$
Lemma 5.
In Algorithm 1, we have that given
$T_{B}\geq 32\left(m\log\left(\frac{192\cdot n\cdot\overline{L}_{h}^{2}}{\underline{L}_{h}^{2}}\right)+\log\left(\frac{T_{G}}{\delta}\right)\right)\sim O\left(m\left(\log(n)+\log\left(\frac{\overline{L}_{h}}{\underline{L}_{h}}\right)\right)+\log\left(\frac{T_{G}}{\delta}\right)\right),$
then with probability at least $1-\delta$
${\lambda_{\min}}(J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top})$ $\geq\frac{\underline{L}_{h}^{2}}{32}-\overline{L}_{h}Rr_{1}^{2}$
$\sigma_{\min}(G_{h})$ $\geq\frac{\underline{L}_{h}^{2}}{32}-\overline{L}_{h}R(r_{1}^{2}+nr_{2}^{2}).$
for all $t=1,2,\dots,T_{G}$
Proof.
From Taylor’s expansion and Assumption 3 we have that
$\widetilde{J}_{h}(x_{t})=\frac{n}{T_{B}}\sum_{i=1}^{T_{b}}J_{h}(x_{t})u_{i}u_{i}^{\top}+\epsilon(x_{t}),$
where $\|\epsilon(x_{t})\|\leq Rr_{1}^{2}$. Thus
$J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}=\frac{n}{T_{B}}\sum_{i=1}^{T_{b}}J_{h}(x_{t})u_{i}u_{i}^{\top}J_{h}(x_{t})+\widetilde{\epsilon}(x_{t}),~~~\textup{where }\|\widetilde{\epsilon}(x_{t})\|\leq\overline{L}_{h}Rr_{1}^{2}$
Further, from Lemma 6 we have that when
$T_{B}\geq 32\left(m\log\left(\frac{192\cdot n\cdot\overline{L}_{h}^{2}}{\underline{L}_{h}^{2}}\right)+\log\left(\frac{T_{G}}{\delta}\right)\right)\sim O\left(m\left(\log(n)+\log\left(\frac{\overline{L}_{h}}{\underline{L}_{h}}\right)\right)+\log\left(\frac{T_{G}}{\delta}\right)\right)$
then with probability at least $1-\delta$
${\lambda_{\min}}\left(\frac{n}{T_{B}}\sum_{i=1}^{T_{B}}J_{h}(x_{t})u_{i}u_{i}^{\top}J_{h}(x_{t})\right)\geq\frac{\underline{L}_{h}^{2}}{32},~~\forall~t=1,2,\dots,T_{G}$
And thus
${\lambda_{\min}}(J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top})\geq\frac{\underline{L}_{h}^{2}}{32}-\overline{L}_{h}Rr_{1}^{2}$
Further, from Taylor expansion, we have that
$\|G_{h}-J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\|\leq n\overline{L}_{h}Rr_{2}^{2}$
and thus
$\sigma_{\min}(G_{h})\geq\frac{\underline{L}_{h}^{2}}{32}-\overline{L}_{h}R(r_{1}^{2}+nr_{2}^{2})$
Proving Lemma 5 will need the following fundamental theorem that uses Small-ball condition to prove anti-concentration:
Theorem 3 (Small-Ball Lower Bound on Minimum Eigenvalue on Empirical Covariance Matrix).
@@ Appendix D Bounding ${\lambda_{\min}}(J_{h}(x_{t})\widetilde{J}_{h}(x_{t}))$
1. (Small-ball condition) For all $z\in\mathbb{S}^{n-1}$:
$\mathbb{P}(|\langle u_{i},z\rangle|\geq\tau)\geq p.$
2. For all $z\in\mathbb{S}^{n-1}$:
$|\langle u_{i},z\rangle|^{2}\leq K.$
Then for any $\delta\in(0,1)$, if:
$N\geq\frac{8}{p^{2}}\left(n\log\left(\frac{24K}{\tau^{2}p}\right)+\log\left(\frac{1}{\delta}\right)\right)$
then with probability at least $1-\delta$,
$\lambda_{\min}\left(\frac{1}{N}\sum_{i=1}^{N}u_{i}u_{i}^{\top}\right)\geq\frac{\tau^{2}p}{4}.$
Proof.
Define:
$S:=\frac{1}{N}\sum_{i=1}^{N}u_{i}u_{i}^{\top}.$
Then for any $z\in\mathbb{S}^{n-1}$:
$z^{\top}Sz=\frac{1}{N}\sum_{i=1}^{N}\langle u_{i},z\rangle^{2}.$
Let $Z_{i}=\langle u_{i},z\rangle^{2}$. By assumption, $\mathbb{P}(Z_{i}\geq\tau^{2})\geq p$. Define indicator variables $A_{i}:=\mathbb{I}\{Z_{i}\geq\tau^{2}\}$. Then $A_{i}\sim\text{Bernoulli}(\geq p)$, and:
$z^{\top}Sz\geq\frac{\tau^{2}}{N}\sum_{i=1}^{N}A_{i}.$
By the Chernoff bound, for all $N\geq 1$,
$\mathbb{P}\left(\sum_{i=1}^{N}A_{i}<\frac{pN}{2}\right)\leq\exp\left(-\frac{p^{2}N}{8}\right).$
Thus, with probability at least $1-\exp\left(-\frac{p^{2}N}{8}\right)$, for a fixed $z$,
$z^{\top}Sz\geq\frac{\tau^{2}p}{2}.$
@@ 2. For all $z\in\mathbb{S}^{n-1}$:
$\mathbb{P}\left(\exists z\in\mathcal{N}_{\epsilon}:z^{\top}Sz<\frac{\tau^{2}p}{2}\right)\leq(3/\epsilon)^{n}\cdot\exp\left(-\frac{p^{2}N}{8}\right).$
To ensure the right hand side is $\leq\delta$, it suffices that:
$N\geq\frac{8}{p^{2}}\left(n\log\left(\frac{3}{\epsilon}\right)+\log\left(\frac{1}{\delta}\right)\right).$
To extend from the net to all $z$, note:
$|z^{\top}Sz-\hat{z}^{\top}S\hat{z}|=\left|(z+\hat{z})S(z-\hat{z})\right|\leq 2K\epsilon$
And thus choosing $\epsilon\sim\frac{\tau^{2}p}{8K}$ ensures for all $z\in\mathbb{S}^{n-1}$:
$z^{\top}Sz\geq\frac{\tau^{2}p}{2}-2K\cdot\epsilon\geq\frac{\tau^{2}p}{4}.$
This concludes the proof. ∎
The following lemma is an immediate corollary of Theorem 3.
Lemma 6.
Given a fixed matrix $A\in\mathbb{R}^{m\times n}$, and random variables $u_{1},u_{2},\dots,u_{N}\in\mathbb{R}^{n}$ where $u_{i}$ is sampled i.i.d. from the unit sphere, we have that when
$N\geq 32\left(m\log\left(192\cdot n\cdot\kappa(AA^{\top})\right)+\log\left(\frac{1}{\delta}\right)\right)\sim O\left(m(\log(n)+\log(\kappa(AA^{\top}))+\log\left(\frac{1}{\delta}\right)\right).$
@@
${\lambda_{\min}}\left(\frac{n}{N}\sum_{i=1}^{N}Au_{i}u_{i}^{\top}A^{\top}\right)\geq\frac{{\lambda_{\min}}(AA^{\top})}{32}$
Proof.
Define $v_{i}=\sqrt{n}Au_{i}$. From the property of random uniform unit sphere vectors Klartag [2017] we have that for any $z\in\mathbb{R}^{n}$,
$\mathbb{P}\left(|u_{i}^{\top}z|\geq\frac{1}{2\sqrt{n}}\|z\|\right)\geq 1/2$
@@
$\mathbb{P}\left(|v_{i}^{\top}z^{\prime}|\geq\frac{1}{2}\sigma_{\min}(A)\|z^{\prime}\|\right)\geq\mathbb{P}\left(|v_{i}^{\top}z^{\prime}|\geq\frac{1}{2}\|Az^{\prime}\|\right)=\mathbb{P}\left(|u_{i}^{\top}z|\geq\frac{1}{2\sqrt{n}}\|z\|\right)\geq 1/2$
Thus, the vector $v_{i}$’s satisfies Condition 1 in Theorem 3 with
$\tau=\frac{\sigma_{\min}(A)}{2},\quad p=\frac{1}{2}.$
Further, given that $u_{i}\in\mathbb{S}^{n-1}$,
$|v_{i}^{\top}z^{\prime}|=\sqrt{n}|u_{i}^{\top}A^{\top}z^{\prime}|\leq\sqrt{n}\|A\|.$
Thus the vector $v_{i}$’s satisfies Condition 2 in Theorem 3 with
$K=n\|A\|^{2}.$
Directly applying Theorem 3 will finish the proof. ∎
Appendix E Auxiliaries
Lemma 7.
$\prod_{\tau=s}^{t-1}(1-\frac{\eta}{\sqrt{\tau+1}})\leq e^{-\eta\left(\sqrt{t}-\sqrt{s+1}\right)}$
Proof.
$\quad\prod_{\tau=s}^{t-1}(1-\frac{\eta}{\sqrt{\tau+1}})\leq\prod_{\tau=s}^{t-1}e^{-\frac{\eta}{\sqrt{\tau+1}}}\leq e^{-\eta\sum_{\tau=s}^{t-1}\frac{1}{\sqrt{\tau+1}}}\leq e^{-\eta(\sqrt{t}-\sqrt{s+1})}$
Lemma 8.
$\sum_{s=1}^{t}e^{\eta\sqrt{s}}\frac{1}{\sqrt{s}}\leq\frac{2}{\eta}2e^{\eta\sqrt{t+1}}$
Proof.
$\sum_{s=1}^{t}e^{\eta\sqrt{s}}\frac{1}{\sqrt{s}}\leq\int_{s=1}^{t+1}e^{\eta\sqrt{s}}\frac{1}{\sqrt{s}}ds=2\int_{s=1}^{t+1}e^{\eta\sqrt{s}}d\sqrt{s}\leq\frac{2}{\eta}e^{\eta\sqrt{t+1}}$
Lemma 9.
$\sum_{s=1}^{t}e^{\eta\sqrt{s}}\frac{1}{s}\leq\frac{e^{2}}{\eta}+\frac{2}{\eta}\frac{1}{\sqrt{t+1}}e^{\eta\sqrt{t+1}}$
Proof.
$\quad\sum_{s=1}^{t}e^{\eta\sqrt{s}}\frac{1}{s}\leq\int_{s=1}^{t+1}e^{\eta\sqrt{s}}\frac{1}{{s}}ds\leq 2\int_{s=1}^{t+1}e^{\eta\sqrt{s}}\frac{1}{\sqrt{s}}d\sqrt{s}$
$=\int_{x=1}^{\sqrt{t+1}}\frac{1}{x}e^{\eta x}dx\leq\int_{x=1}^{\frac{2}{\eta}}e^{\eta x}dx+\int_{x=\frac{2}{\eta}}^{\sqrt{t+1}}\frac{1}{x}e^{\eta x}dx$
$\leq\int_{x=1}^{\frac{2}{\eta}}e^{\eta x}dx+\int_{x=\frac{2}{\eta}}^{\sqrt{t+1}}\left(\frac{2}{x}-\frac{2}{\eta x^{2}}\right)e^{\eta x}dx\leq\frac{e^{2}}{\eta}+\frac{2}{\eta}\frac{1}{\sqrt{t+1}}e^{\eta\sqrt{t+1}}$

### v2

@@ Appendix E Bounding ${\lambda_{\min}}(J_{h}(x_{t})\widetilde{J}_{h}(x_{t}))$
$\textstyle{\lambda_{\min}}(J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top})\geq\frac{\underline{L}_{h}^{2}}{32}-\overline{L}_{h}Rr_{1}^{2}$
Further, from Lemma F.1 we have $\|G_{h}-J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\|\leq n\overline{L}_{h}Rr_{2}^{2}$ and thus
$\textstyle\sigma_{\min}(G_{h})\geq\frac{\underline{L}_{h}^{2}}{32}-\overline{L}_{h}R(r_{1}^{2}+nr_{2}^{2})$
Proving Lemma E.1 will need the following fundamental theorem that uses Small-ball condition to prove anti-concentration:
Theorem E.3 (Small-Ball Lower Bound on Minimum Eigenvalue on Empirical Covariance Matrix).
@@
  1. 1.
    (Small-ball condition) For all $z\in\mathbb{S}^{n-1}$: $\mathbb{P}(|\langle u_{i},z\rangle|\geq\tau)\geq p.$
  2. 2.
    For all $z\in\mathbb{S}^{n-1}$: $|\langle u_{i},z\rangle|^{2}\leq K.$
Then for any $\delta\in(0,1)$, if $N\geq\frac{8}{p^{2}}\left(n\log\left(\frac{24K}{\tau^{2}p}\right)+\log\left(\frac{1}{\delta}\right)\right)$ then with probability at least $1-\delta$, $\lambda_{\min}\left(\frac{1}{N}\sum_{i=1}^{N}u_{i}u_{i}^{\top}\right)\geq\frac{\tau^{2}p}{4}.$
Proof E.4.
Define $S:=\frac{1}{N}\sum_{i=1}^{N}u_{i}u_{i}^{\top}.$ Then for any $z\in\mathbb{S}^{n-1}$ we have $z^{\top}Sz=\frac{1}{N}\sum_{i=1}^{N}\langle u_{i},z\rangle^{2}$ Let $Z_{i}=\langle u_{i},z\rangle^{2}$. By assumption, $\mathbb{P}(Z_{i}\geq\tau^{2})\geq p$. Define indicator variables $A_{i}:=\mathbb{I}\{Z_{i}\geq\tau^{2}\}$. Then $A_{i}\sim\text{Bernoulli}(\geq p)$, and $z^{\top}Sz\geq\frac{\tau^{2}}{N}\sum_{i=1}^{N}A_{i}.$ By the Chernoff bound, for all $N\geq 1$, $\mathbb{P}\left(\sum_{i=1}^{N}A_{i}<\frac{pN}{2}\right)\leq\exp\left(-\frac{p^{2}N}{8}\right).$ Thus, with probability at least $1-\exp\left(-\frac{p^{2}N}{8}\right)$, for a fixed $z$, $z^{\top}Sz\geq\frac{\tau^{2}p}{2}.$
@@
$\textstyle\mathbb{P}\left(\exists z\in\mathcal{N}_{\epsilon}:z^{\top}Sz<\frac{\tau^{2}p}{2}\right)\leq(3/\epsilon)^{n}\cdot\exp\left(-\frac{p^{2}N}{8}\right).$
To ensure the right hand side is $\leq\delta$, it suffices that: $N\geq\frac{8}{p^{2}}\left(n\log\left(\frac{3}{\epsilon}\right)+\log\left(\frac{1}{\delta}\right)\right).$ To extend from the net to all $z$, note:
$\textstyle|z^{\top}Sz-\hat{z}^{\top}S\hat{z}|=\left|(z+\hat{z})S(z-\hat{z})\right|\leq 2K\epsilon$
And thus choosing $\epsilon\sim\frac{\tau^{2}p}{8K}$ ensures for all $z\in\mathbb{S}^{n-1}$: $z^{\top}Sz\geq\frac{\tau^{2}p}{2}-2K\cdot\epsilon\geq\frac{\tau^{2}p}{4}.$ This concludes the proof.
The following lemma is an immediate corollary of Theorem E.3.
Lemma E.5.
Given a fixed matrix $A\!\in\!\mathbb{R}^{m\times n}$ and random variables $u_{1},\!u_{2},\dots,\!u_{N}\!\in\!\mathbb{R}^{n}$ where $u_{i}$ is sampled i.i.d. from the unit sphere, we have that when
[…]
@@
$\textstyle{\lambda_{\min}}\left(\frac{n}{N}\sum_{i=1}^{N}Au_{i}u_{i}^{\top}A^{\top}\right)\geq\frac{{\lambda_{\min}}(AA^{\top})}{32}$
Proof E.6.
Define $v_{i}=\sqrt{n}Au_{i}$. From the property of random uniform unit sphere vectors [klartag2017super] we have that for any $z\in\mathbb{R}^{n}$,
$\textstyle\mathbb{P}\left(|u_{i}^{\top}z|\geq\frac{1}{2\sqrt{n}}\|z\|\right)\geq 1/2$
@@
$\textstyle\mathbb{P}\left(|v_{i}^{\top}z^{\prime}|\geq\frac{1}{2}\sigma_{\min}(A)\|z^{\prime}\|\right)\geq\mathbb{P}\left(|v_{i}^{\top}z^{\prime}|\geq\frac{1}{2}\|Az^{\prime}\|\right)=\mathbb{P}\left(|u_{i}^{\top}z|\geq\frac{1}{2\sqrt{n}}\|z\|\right)\geq 1/2$
Thus, the vector $v_{i}$’s satisfies Condition 1 in Theorem E.3 with $\tau=\frac{\sigma_{\min}(A)}{2},\quad p=\frac{1}{2}.$ Further, given that $u_{i}\in\mathbb{S}^{n-1}$, $|v_{i}^{\top}z^{\prime}|=\sqrt{n}|u_{i}^{\top}A^{\top}z^{\prime}|\leq\sqrt{n}\|A\|.$ Thus the vector $v_{i}$’s satisfies Condition 2 in Theorem E.3 with $K=n\|A\|^{2}.$ Directly applying Theorem E.3 will finish the proof.
Appendix F Auxiliaries
Lemma F.1.
For $\widetilde{J}(x)t),\widetilde{\nabla}f(x_{t})$ defined in (7) and $G_{h},G_{f}$ in (11), we have
$\textstyle\|G_{h}\!-\!J_{h}(x_{t})\widetilde{J}_{h}(x_{t})^{\top}\|\!\leq\!n\overline{L}_{h}Rr_{2}^{2},~\|G_{f}\!-\!J_{h}(x_{t})\widetilde{\nabla}f(x_{t})\|\!\leq\!nL_{f}Rr_{2}^{2},~\|\widetilde{\nabla}f(x_{t})\|\!\leq\!nL_{f}.$
Proof F.2.
The first two inequalities obtained from standard truncation-error bounds for the central difference directional derivative [suli2003introduction]. The last inequality can be derived by the Lagrange’s Mean Value Theorem, where we can show that for any unit vector $u$, $\|\frac{f(x+ru)-f(x-ru)}{2r}u\|=\|uu^{\top}\nabla f(x+\bar{r}u)\|\leq L_{f}$.
Lemma F.3.
For any $\eta\!>\!0,t,s\!\in\!\mathbb{N},t\!\geq\!s\!\geq\!0$: $\prod_{\tau=s}^{t-1}(1-\frac{\eta}{\sqrt{\tau+1}})\leq e^{-\eta\left(\sqrt{t}-\sqrt{s+1}\right)}.$
Proof F.4.
$\quad\prod_{\tau=s}^{t-1}(1-\frac{\eta}{\sqrt{\tau+1}})\leq\prod_{\tau=s}^{t-1}e^{-\frac{\eta}{\sqrt{\tau+1}}}\leq e^{-\eta\sum_{\tau=s}^{t-1}\frac{1}{\sqrt{\tau+1}}}\leq e^{-\eta(\sqrt{t}-\sqrt{s+1})}$.
Lemma F.5.
For any $\eta>0$ and $t,s\in\mathbb{N},t\geq s\geq 0$: $\sum_{s=1}^{t}e^{\eta\sqrt{s}}\frac{1}{\sqrt{s}}\leq\frac{2}{\eta}2e^{\eta\sqrt{t+1}}$
Proof F.6.
$\sum_{s=1}^{t}e^{\eta\sqrt{s}}\frac{1}{\sqrt{s}}\leq\int_{s=1}^{t+1}e^{\eta\sqrt{s}}\frac{1}{\sqrt{s}}ds=2\int_{s=1}^{t+1}e^{\eta\sqrt{s}}d\sqrt{s}\leq\frac{2}{\eta}e^{\eta\sqrt{t+1}}$.
Lemma F.7.
For any $\eta>0$ and $t,s\in\mathbb{N},t\geq s\geq 0$
$\textstyle\sum_{s=1}^{t}e^{\eta\sqrt{s}}\frac{1}{s}\leq\frac{e^{2}}{\eta}+\frac{2}{\eta}\frac{1}{\sqrt{t+1}}e^{\eta\sqrt{t+1}}$
Proof F.8.
$\textstyle\quad\sum_{s=1}^{t}e^{\eta\sqrt{s}}\frac{1}{s}\leq\int_{s=1}^{t+1}e^{\eta\sqrt{s}}\frac{1}{{s}}ds\leq 2\int_{s=1}^{t+1}e^{\eta\sqrt{s}}\frac{1}{\sqrt{s}}d\sqrt{s}$
$\textstyle=\int_{x=1}^{\sqrt{t+1}}\frac{1}{x}e^{\eta x}dx\leq\int_{x=1}^{\frac{2}{\eta}}e^{\eta x}dx+\int_{x=\frac{2}{\eta}}^{\sqrt{t+1}}\frac{1}{x}e^{\eta x}dx$
$\textstyle\leq\int_{x=1}^{\frac{2}{\eta}}e^{\eta x}dx+\int_{x=\frac{2}{\eta}}^{\sqrt{t+1}}\left(\frac{2}{x}-\frac{2}{\eta x^{2}}\right)e^{\eta x}dx\leq\frac{e^{2}}{\eta}+\frac{2}{\eta}\frac{1}{\sqrt{t+1}}e^{\eta\sqrt{t+1}}.$